4. Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b
mx + y = 22

Answer by fractalier(6550)

We have to plug in -3 for x and 4 for y...we get
5(-3) + 7(4) = -15 + 28 = 13 = b
and then
-3m + 4 = 22
-3m = 18
m = -6